OpenStudy (anonymous):

A snail can move "i" inches in "m" minutes. At this rate, how many feet can it move in "h" hours? 5hi/m 60hi/m hi/12m 5m/hi 5him

OpenStudy (anonymous):

how many minutes are there in an hour?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

right! if the snail can move i inches in "m" minutes. how many inches can it move in 60 minutes?what mathematical operation we use here?

OpenStudy (unklerhaukus):

\[1m=1m\times\frac1{ 60}\frac {h}m=\frac 1{60}h\] \[1i=1i \times\frac1{ 12}\frac {ft}i=\frac 1{12}ft\] \[\frac im=\frac{1/12~ ft}{1/60~ h}\] \[\frac{12}{60}\frac im=\frac{ft}{h}\]